A net force of 1.6×10−15 N acts on an electron over a displacement of 5.0 cm, in the same direction as the net force. (a) What is the change in kinetic energy of the electron? (b) If the electron was initially at rest, what is the speed of the electron? An electron has a mass of 9.1×10−31 kg. answer part a) in joules and part b) in meters per second.
1
Expert's answer
2011-03-09T06:46:08-0500
a) the change in kinetic energy equals to the work done by the force (if the potential energy remains the same):
A = F s = 1.6×10−15[N] * 0.05 [m] = 0.8×10−16J.

b) if the electron was initially at rest, the speed of the electron can be found from the expression
A = mv2/2: v = √(2A/m) = √(2*0.8×10−16J/9.1×10−31) = 0.17×1015m/s